OpenStudy (anonymous):

At the end of the winter, all skis at a sporting goods store were on sale for 3/5 off. What was the sale price of a snowboard that originally cost $360? Answer $120 $240 $216 $144

OpenStudy (anonymous):

@matricked help

OpenStudy (anonymous):

@privetek

OpenStudy (anonymous):

(1-3/5)*360 =2/5 *360 =144